p(x) = k x (x – 1), x = 1, 2, 3, 4.
Obtain the value of k
Construct the probability distribution table for x
Hence, calculate E(x), E(x2) and Var(x)
1
Expert's answer
2019-08-20T12:02:57-0400
"\sum p(x_i)=1;
\\k(1*(1-1)+2*(2-1)+3*(3-1)+
\\+4*(4-1))=1;
\\k=0.05;"

"\begin{matrix}
 x & 1 & 2 & 3 & 4 \\
 p(x) & 0 & 0.1 & 0.3 & 0.6
\end{matrix}"

"E(x)=\sum p(x_i)x_i;
\\E(x)=0*1+0.1*2+0.3*3+0.6*4=3.5;"

"E(x^2)=\sum p(x_i)x_i^2;
\\E(x^2)=0*1+0.1*4+0.3*9+0.6*16=12.7;"

"Var(x)=E(x^2)-E(x)^2;
\\Var(x)=12.7-12.25=0.45;"